1.10) The average weight of 3-year-old boys is 31.5 pounds with a standard deviation of 2.1. Henry is a 3-year-old who weighs 30 pounds. What is the z score for Henry’s weight?

a.

.71

b.

-.71

c.

.5

d.

-.5

1.16)

Two teachers have a friendly rivalry. They want to know which of their students did better on their respective final exams. The first teacher teaches chemistry and her 49 students had a mean = 68. The national mean and standard deviation for this exam is 65 and 10, respectively. The second teacher teaches history and his 64 students had a mean = 72. The national mean and standard deviation for this exam is 70 and 7, respectively. Which teachers’ students did better on their respective final exams?

a.

the chemistry class did better

b.

the history class did better

c.

the two classes did equally as well
Answer #1

Answer 1.10

The Z score is given by:

Z = \frac{X-\mu}{\sigma} = \frac{30 - 31.5}{2.1} = -0.714285714 \sim -0.71

Hence the answer is :-

b.

-.71

Answer 1.16

We calculate the z scores for both students in order to facilitate comparison:

Chemistry

Z = \frac{\bar X-\mu}{\sigma/\sqrt n} =\frac{68-65}{10/\sqrt 49} = 2.1

History

Z = \frac{\bar X-\mu}{\sigma/\sqrt n} =\frac{72-70}{7/\sqrt 64} = 2.285714286

History teacher's z score is higher, hence she did better

b.

the history class did better

Please do upvote if you are satisfied! Let me know in the comments if anything is not clear. I will reply ASAP!